OpenStudy (anonymous):

If Argon effuses at a rate of 1.6 *10^-3 mol in 215 seconds. How much O2 would effuse under the same conditions in the same amount of time?

OpenStudy (jfraser):

use graham's law of effusion: the rates of effusion of two gases is inversely proportional to the square root of their molar masses\[\frac{rate_1}{rate_2} = \sqrt{\frac{MM_2}{MM_1}}\]